# High-dimensional sliced inverse regression with endogeneity

### v1

@@ Acknowledgments
$\left\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\right\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\bm{\Sigma}_{\widehat{\mathbf{X}}}\hat{\bm{\gamma}}_{k}|\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}.$
@@
$\text{trace}(\bm{\eta}^{\top}\hat{\bm{\Lambda}}\bm{\eta})\leq\text{trace}(\hat{\bm{\eta}}^{\top}\hat{\bm{\Lambda}}\hat{\bm{\eta}}),$
@@
$\langle\bm{\Lambda},\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle$ $=\text{trace}(\bm{\Lambda})-\text{trace}(\bm{\eta}\text{diag}(\lambda_{1},\ldots,\lambda_{d})\bm{\eta}^{\top}\widehat{\bm{\eta}}\widehat{\bm{\eta}}^{\top})$
$=\text{trace}(\bm{\Lambda})-\text{trace}(\text{diag}(\lambda_{1},\ldots,\lambda_{d})\mathbf{H}^{2})$
$=\sum_{k=1}^{d}\lambda_{k}(1-\mathbf{H}^{2}_{kk})\geq\lambda_{d}(d-\text{trace}(\mathbf{H}^{2})).$
@@
$\langle\hat{\bm{\Lambda}}-\bm{\Lambda},\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle$ $\leq\left\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max}\left\lVert\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\right\rVert_{1}$
$\leq\sqrt{d}\left\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max}$
$\leq CL\sqrt{d}\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$
@@
$\lambda_{d}\left\{d-\text{trace}(\mathbf{H}^{2})\right\}\leq CL\sqrt{d}\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$
@@
S4 Additional simulation results
Table S4: Simulation results for the Lasso SIR (LSIR) estimator, the two-stage Lasso (2SLasso) estimator, and the proposed two-stage Lasso SIR (2SLSIR) estimator, in the single index models (i) and (ii) and the instrumental variables $\mathbf{Z}$ are independently generated from the signed Bernoulli distribution. Standard errors are included in parentheses. The lowest estimation error and the highest F1 score are highlighted in each row.
$n$ Model $\bm{\Gamma}$ Error F1
LSIR 2SLasso 2SLSIR LSIR 2SLasso 2SLSIR
500 (i) Strong 0.28 (0.07) 1.05 (0.22) 0.13 (0.04) 0.99 (0.03) 0.52 (0.18) 1.00 (0.00)
Weak 0.37 (0.08) 1.01 (0.24) 0.18 (0.06) 0.96 (0.07) 0.54 (0.18) 0.99 (0.03)
Mixed 0.29 (0.08) 1.04 (0.20) 0.11 (0.03) 0.98 (0.05) 0.57 (0.18) 1.00 (0.00)
(ii) Strong 0.26 (0.08) 0.24 (0.07) 0.13 (0.05) 1.00 (0.02) 0.98 (0.04) 1.00 (0.00)
Weak 0.35 (0.07) 0.31 (0.08) 0.18 (0.07) 0.97 (0.05) 0.94 (0.08) 1.00 (0.01)
Mixed 0.30 (0.06) 0.24 (0.08) 0.11 (0.04) 0.99 (0.03) 0.98 (0.05) 1.00 (0.00)
1000 (i) Strong 0.30 (0.07) 0.98 (0.22) 0.09 (0.03) 0.99 (0.03) 0.57 (0.18) 1.00 (0.00)
Weak 0.38 (0.07) 0.90 (0.23) 0.13 (0.04) 0.96 (0.06) 0.63 (0.16) 1.00 (0.01)
Mixed 0.35 (0.06) 0.89 (0.24) 0.06 (0.02) 0.97 (0.07) 0.63 (0.17) 1.00 (0.00)
(ii) Strong 0.31 (0.07) 0.18 (0.05) 0.09 (0.03) 0.99 (0.04) 1.00 (0.02) 1.00 (0.00)
Weak 0.39 (0.07) 0.22 (0.07) 0.12 (0.04) 0.94 (0.08) 0.98 (0.05) 1.00 (0.01)
Mixed 0.34 (0.06) 0.16 (0.05) 0.07 (0.02) 0.97 (0.07) 1.00 (0.01) 1.00 (0.00)
Table S5: Simulation results for the Lasso SIR (LSIR) estimator, the two-stage Lasso (2SLasso) estimator, and the proposed two-stage Lasso SIR (2SLSIR) estimator, in double index models (iii) and (iv) and the instrumental variables $\mathbf{Z}$ are independently generated from the signed Bernoulli distribution. Standard errors are included in parentheses. The lowest estimation error and the highest F1 score are highlighted in each row.
$n$ Model $\bm{\Gamma}$ Error F1
LSIR 2SLSIR LSIR 2SLSIR
500 (iii) Strong 0.88 (0.23) 0.67 (0.26) 0.61 (0.16) 0.73 (0.16)
Weak 1.02 (0.21) 0.92 (0.29) 0.56 (0.12) 0.62 (0.17)
Mixed 0.98 (0.19) 0.70 (0.26) 0.49 (0.11) 0.70 (0.16)
(iv) Strong 0.40 (0.21) 0.40 (0.24) 0.93 (0.11) 0.93 (0.13)
Weak 0.50 (0.20) 0.49 (0.25) 0.88 (0.13) 0.85 (0.15)
Mixed 0.43 (0.21) 0.39 (0.23) 0.88 (0.14) 0.91 (0.14)
1000 (iii) Strong 0.97 (0.21) 0.49 (0.29) 0.55 (0.13) 0.86 (0.15)
Weak 1.08 (0.20) 0.66 (0.29) 0.53 (0.10) 0.73 (0.18)
Mixed 1.05 (0.16) 0.50 (0.28) 0.45 (0.07) 0.84 (0.14)
(iv) Strong 0.38 (0.21) 0.31 (0.24) 0.95 (0.13) 0.97 (0.11)
Weak 0.49 (0.22) 0.35 (0.24) 0.88 (0.15) 0.93 (0.13)
Mixed 0.43 (0.22) 0.29 (0.22) 0.88 (0.16) 0.96 (0.12)
S5 Additional results for the application
Figure S4: Heatmap corresponding to the absolute value of the elements of $\widehat{\bm{\Gamma}}$ from the first stage of the 2SLSIR estimator applied to the NHANES data. Each column represents the consumption of one dietary component in the first interview, which is treated as a covariate. Each column represents the consumption of one dietary component in the second interview, which is treated as an instrument.

### v2

@@ C.3 Proof of Lemma 38
$\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\bm{\Sigma}_{\widehat{\mathbf{X}}}\hat{\bm{\gamma}}_{k}|\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}.$
@@ C.4 Proof of Proposition 5
or equivalently $\langle\hat{\bm{\Lambda}},\bm{\eta}\bm{\eta}^{\top}\rangle\leq\langle\hat{\bm{\Lambda}},\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle$, where for two matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ we let $\langle\mathbf{A}_{1},\mathbf{A}_{2}\rangle=\text{trace}(\mathbf{A}_{1}^{\top}\mathbf{A}_{2})$ denote their dot product. Writing $\hat{\bm{\Lambda}}=\bm{\Lambda}+(\hat{\bm{\Lambda}}-\bm{\Lambda})$, it follows that
@@
$\langle\bm{\Lambda},\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle$ $=\text{trace}(\bm{\Lambda})-\text{trace}\{\bm{\eta}\text{diag}(\lambda_{1},\ldots,\lambda_{d})\bm{\eta}^{\top}\widehat{\bm{\eta}}\widehat{\bm{\eta}}^{\top}\}$
$=\text{trace}(\bm{\Lambda})-\text{trace}\{\text{diag}(\lambda_{1},\ldots,\lambda_{d})\mathbf{H}^{2}\}$
$=\sum_{k=1}^{d}\lambda_{k}(1-h^{2}_{kk})\geq\lambda_{d}(d-\text{trace}(\mathbf{H}^{2}).$
@@
$\langle\hat{\bm{\Lambda}}-\bm{\Lambda},\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle$ $\leq\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\rVert_{\max}\lVert\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rVert_{1}$
$\leq\sqrt{d}\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\rVert_{\max}$
$\leq CL\sqrt{d}\left(L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right),$
@@
$\lambda_{d}\left\{d-\text{trace}(\mathbf{H}^{2})\right\}\leq CL\sqrt{d}\left(L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right),$
@@
$\text{trace}(\mathbf{H}^{2})\geq d-\dfrac{CL\sqrt{d}}{\lambda_{d}}\left(L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right),$
@@
Appendix D Additional results for the application
D.1 Mouse obesity data
In this subsection, we further interpret the biological consequences of the genes selected by different estimators as outlined in Section 7.1 of the main paper. To that end, we conducted a gene ontology enrichment analysis for the genes chosen by each method. By annotating these genes for their participation in gene ontology, this analysis aims to identify which biological functions are over-represented in the selected genes (Bleazard et al.,, 2015). This over-representation is measured by the $p$-values, adjusted for multiple comparisons, of the null hypothesis that this gene set is randomly associated with the GO. A smaller $p$-value thus indicates a more significant biological function. Furthermore, the gene ontology is divided into three main groups: biological processes (BP), molecular function (MF), and cellular component (CC). Figure 3 indicates that all the three methods capture a strong association between the obesity and lipid-related processes, including lipid transport, metabolism, and regulation. The one-stage lasso SIR estimator focuses on genes regulating lipid transport and metabolic processes (e.g., lipid transporter activity, monocarboxylic acid transport).
By contrast, both the two stage lasso and two-stage lasso SIR estimators emphasize plasma lipoprotein metabolism (e.g., plasma lipoprotein particle clearance, protein-lipid complex), with the former putting more emphasis on the interplay between lipid metabolism and immune regulation (e.g., immunoglobulin binding) while the latter focuses more on lipid catabolism and extracellular lipid transport (e.g., extracellular matrix structural constituent, lipase activity).
D.2 NHANES data
Figure 4: Heatmap corresponding to the absolute value of the elements of $\widehat{\bm{\Gamma}}$ from the first stage of the 2SLSIR estimator applied to the NHANES data. Each column represents the consumption of one dietary component in the first interview, which is treated as a covariate. Each column represents the consumption of one dietary component in the second interview, which is treated as an instrument.